OpenStudy (anonymous):

what is the slop between the points (1,5) and (5,2)

OpenStudy (anonymous):

from 1 to 5 is right 4 units, from 5 to 2 is down 3 right 4, down 3 slope is \(-\frac{3}{4}\)

OpenStudy (anonymous):

you can also use \[\frac{y_2-y_1}{x_2-x_1}=\frac{2-5}{5-1}\] you will get the same answer